OpenStudy (anonymous):

MEDAL TO WHOEVER ANSWERS (5 questions) Unit 2 test review, due today. Tell whether the given number is a solution of each equation. 4. 5-3x=11;x=-2 5.x/2-5=3;x=16 Tell wheather the orderd pair is a solution to the given number 6. y=-2 ; (5,-10) 7. y=3x+9 ; (1,12) 8.Charlie is 4 years older than his sister Sally. Let x represent Charlies age and y represent Sallys age. Create an equation and a graph to represent the relationship

OpenStudy (anonymous):

\[ 5-3x=11;x=-2\] replace \(x\) by \(-2\) and see if you get \(11\), i.e. compute \[2-3\times (-2)\] if it is \(11\) then "yes" if it is not \(11\) then "no"

OpenStudy (studygurl14):

this is easy. All you do is plug in the num,bers and see if they work. (4-7) If they don't then they are not a solution, if they do, then they are. For 8, x = Charlie y= sally x = y + 4 --> y = -x + 4

OpenStudy (anonymous):

hello \(\large \begin{array}l\color{red}{\normalsize\text{S}}\color{orange}{\normalsize\text{A}}\color{#9c9a2e}{\normalsize\text{I}}\color{green}{\normalsize\text{F}}\color{blue}{\normalsize\text{O}}\color{purple}{\normalsize\text{O}}\color{purple}{\normalsize\text{O}}\color{red}{\normalsize\text{O}}\color{orange}{\normalsize\text{O}}\color{#9c9a2e}{\normalsize\text{O}}\color{green}{\normalsize\text{!}}\color{blue}{\normalsize\text{!}}\color{purple}{\normalsize\text{}}\end{array}\)

OpenStudy (anonymous):

thanks i only understand 8 though. if yu can answer the rest, thatd be pretty cool

OpenStudy (studygurl14):

Okay @prettyfeather here is the solution to #4. Hopefully it helps with the rest. Tell whether the given number is a solution of each equation. 4. 5-3x=11;x=-2 Plug in -2 into the equation as x to find whether it makes the equation true. 5 - 3x = 11 5 - 3(-2) =11 5 + 6 = 11 True So, yes, x=-2 is a solution of the equation. Hope this helps. :)

OpenStudy (studygurl14):

hm...don't know why it typed "feather". I put @prettyfeather

OpenStudy (studygurl14):

odd...it did it again...